✓ Saved Question 6 (1 point) What is the y-intercept of y − 3x – 12 = 0? -12 3 12 -3
Answer #1

Salmi given that y - 3X - 12 20 y = 3x +12 Put x=0, in eah o, we get, 3x0 +12 = y-intercept = 12 Ars:- (c)